OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3).

OpenStudy (anonymous):

Equation of the required line is given by: \[\large \frac{y - y_1}{y _ 2 - y_1} = \frac{x - x_1}{x _ 2 - x_1}\]

OpenStudy (anonymous):

Plug in the values: \(x_1 = 8\) \(y_1 = 3\) \(x_2 = 4\) \(y_2 = 3\)

OpenStudy (anonymous):

Or you can multiply both the sides by \(y_2 - y_1\) \[\large y - y_1= (\frac{x - x_1}{x _ 2 - x_1}) \times (y_2 - y_1)\]

OpenStudy (anonymous):

I can't get past finding the slope.

OpenStudy (anonymous):

I have a misunderstanding there.

OpenStudy (anonymous):

So: \[y - 3 = \frac{x - 8}{4-8} \times (3-3)\] \(y - 3 = 0\) \(y = 3\)

OpenStudy (anonymous):

Or you can find slope first.. I think you got slope as 0.. So use: \[\large y-y_1 = m(x - x_1)\] Here m is slope that is 0. so, \(y - y_1 = 0(x - x_1)\) \(y - y_1 = 0\) \(y_1\) is 3 here: \[y-3 = 0\] \(y=3\)